how many liters of a 20% alcohol solution must be mixed with 30 liters of a 80% solution to get a 40% solution?
Answer by robertb(5830)

0.20x + 0.80*30 = 0.40(x + 30),
0.20x + 24 = 0.40x + 12,
12 = 0.20x,
x = 60, #liters of 20% alcohol solution that must be added.